OpenStudy (jenniferjuice):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

When they ask for something like this you just place g(x) into all x for f(x) so it looks like f(g(x))= -4(10x-6) + 7 you can simplify this by expanding the bracket.

OpenStudy (jenniferjuice):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (jenniferjuice):

i got -9x + 29

OpenStudy (jenniferjuice):

but none of my choices are that

OpenStudy (marissalovescats):

Well plug g(x) in the x for f(x) 9(-x+3)-2 -9x+27-2 -9x+25

OpenStudy (marissalovescats):

You added 2 instead of subtracted it

OpenStudy (jenniferjuice):

ohhh okay thanks

OpenStudy (marissalovescats):

You're welcome